In a lab one student is forming an experiment in which he mixes 2L of 0.1M H2SO4
at 298 K with 4L of 0.1M NaOH also at 298 K. What will be the final temperature
after this mixing, if the specific heat capacity and density of final mixture are 0.9 J/g-
0C and 1.05 g/ml respectvely
1
Expert's answer
2021-07-12T03:00:09-0400

P1V1/n1T1= P2V2/n2T2

PV = nRT.

2/298×0.1 = 4/298×0.1

0.067 = 0.13423

= (0.067 + 0.1323) × 0.9×1.05× 273

= 51.42K